OpenStudy (anonymous):

3x^2+2x=0 quadratic equation

OpenStudy (anonymous):

solve?

OpenStudy (anonymous):

factor the left side.... what do you have....??

OpenStudy (anonymous):

solve

OpenStudy (anonymous):

what's the greatest common factor on the left side?

OpenStudy (anonymous):

show steps

OpenStudy (anonymous):

ok... i'll show steps but you need to do this with me and answer the question: what's the greatest common factor on the left side?

OpenStudy (anonymous):

i dont know how to do it

OpenStudy (anonymous):

on the left side of the equation you have two terms: \(\large 3x^2=3\cdot x \cdot x \) and \(\large 2x = 2 \cdot x \) what is the factor they both have in common?